# TimeBridge: Non-Stationarity Matters for Long-term Time Series Forecasting

### v2

@@ 1 Introduction
Recently, many methods have emerged to utilize a normalization-and-denormalization paradigm to address non-stationarity in time series (Kim et al., 2022; Fan et al., 2023; Liu et al., 2023; 2024b). For instance, RevIN (Kim et al., 2022) normalizes the input data and subsequently applies its distributional characteristics to denormalize the output predictions. Building on this approach, other methods have designed distributional prediction networks (Fan et al., 2023) and more refined normalization techniques (Liu et al., 2023) to further mitigate non-stationarity. On the other hand, some studies (Liu et al., 2022b; Ma et al., 2024; Fan et al., 2024) argue that over-stabilizing time series may actually reduce the richness of embedded features, leading to a decline in model performance. Existing methods for addressing non-stationarity in time series face a dilemma: some prioritize eliminating non-stationary factors to reduce overfitting, while others attempt to incorporate these factors but lack comprehensive theoretical frameworks. Furthermore, they do not adequately explain the trade-offs between removing non-stationarity and leveraging it for modeling.
It is notable that non-stationary characteristics have distinct impacts on modeling short-term and long-term dependencies. Non-stationarity can lead to spurious regressions in short-term modeling due to the high randomness and unpredictability of short-term fluctuations (Noriega & Ventosa-Santaulària, 2007) (see Fig. 1b). For example, a sudden drop in temperature could be caused by a typhoon or cold front, both of which have no intrinsic connection. Retaining non-stationarity can result in false correlations between such unrelated events. However, non-stationarity is crucial for capturing long-term cointegration relationships among variates, reflecting their co-movement or synchronized changes over time (Fanchon & Wendel, 1992). Removing non-stationarity may also eliminate these essential long-term dependencies (see Fig. 1c). Therefore, while non-stationarity can cause spurious regressions in short-term modeling, it is essential for modeling long-term dependencies between variates. Conversely, eliminating non-stationarity benefits short-term modeling but erases long-term cointegration (see Fig. 1d).
In this paper, to address the dual challenges posed by non-stationarity in short-term and long-term modeling, we propose distinct strategies tailored to each scenario. For short-term modeling, we eliminate non-stationarity to capture the strong temporal dependencies within each variate, as short-term causal relationships exist mainly between consecutive time points within a single variate rather than across variates. This strategy reduces the risk of spurious regressions from non-stationary fluctuations, enabling the model to better capture the local causal dynamics. For long-term modeling, we utilize the preserved non-stationarity to uncover long-term cointegration relationships between different variates, thereby enabling more accurate and reliable long-term forecasting.
@@
Figure 1: Visualization of the impact of non-stationarity on short-term and long-term modeling. (a) Comparison of two cointegrated series $X_{t}$ and $Y_{t}$ before and after de-trending, with $X_{t}$ showing two random fluctuations in Phase A. (b) Short-term patch attention map of $X_{t}$ in Phase A, showing that non-stationarity leads to spurious regression in short-term modeling. (c) Long-term relationship between $X_{t}$ and $Y_{t}$, where removing non-stationarity eliminates the cointegration. (d) Non-stationarity enables modeling long-term cointegration between variates but causes spurious regressions in short-term modeling (orange line), while de-trending benefits short-term modeling but disrupts long-term relationships (green line).
@@ 2 Related Works
As shown in Fig. 2, recent advancements in multivariate time series forecasting have predominantly focused on two core directions: Normalization and Dependecy Modeling.
Normalization can be divided into stationary and non-stationary methods. Stationary methods (Kim et al., 2022; Fan et al., 2023; Liu et al., 2024b; 2023) aim to eliminate non-stationarity through model-agnostic normalization techniques, thereby preventing spurious regressions and enhancing model performance. For example, RevIN (Kim et al., 2022) applies Z-normalization to the input sequence and then reverses the normalization on the output using the distributional characteristics of the input, assuming that both share similar distributional properties. Dish-TS (Fan et al., 2023) takes this further by predicting the statistical characteristics of the output with a distribution prediction model. Additionally, SAN (Liu et al., 2023) offers a more granular patch-level prediction method. Conversely, some approaches (Liu et al., 2022b; Ma et al., 2024; Fan et al., 2024) advocate preserving non-stationarity, as excessive normalization can eliminate diverse sequence characteristics and limit predictive accuracy.
@@ 3.1 Structure Overview
As illustrated in Fig. 3, our proposed TimeBridge consists of four key components: (a) Patch Embedding segments the input sequence into non-overlapping patches and transforms each patch into a patch token; (b) Integrated Attention models the dependencies among all patch tokens of the same variates. By eliminating non-stationarity within each patch token, it mitigates the risk of spurious regressions that could arise from abrupt short-term changes; (c) Patch Downsampling aggregates global information and reduces the number of patches to encapsulate richer long-term features within each patch, while simultaneously lowering computational complexity; (d) Cointegrated Attention preserves the non-stationary characteristics of the sequence and models the long-term cointegration relationships across different variates within the same temporal window.
@@ 3.2 Patch Embedding
$\{\mathbf{p}_{1},\cdots,\mathbf{p}_{N}\}=\text{Patching}(\mathbf{X}),\quad\mathbf{P}=\text{Embedding}(\mathbf{p}_{1},\cdots,\mathbf{p}_{N})$ (1)
@@ 3.3 Integrated Attention
$\mathbf{p}_{i}^{\text{Trend}}=\text{AvgPool}(\text{Padding}(\mathbf{p}_{i})),\quad\mathbf{p}_{i}^{\prime}=\mathbf{p}_{i}-\mathbf{p}_{i}^{\text{Trend}},\quad\mathbf{P}_{c,:}^{\prime}=\{\mathbf{p}_{1}^{\prime},\cdots,\mathbf{p}_{N}^{\prime}\},$ (2)
@@ 3.4 Patch Downsampling
$\mathbf{P}_{c,:}^{\prime}=\text{Downsample}(\mathbf{P}_{c,:}),\quad\mathbf{P}_{c,:}=\text{Attention}(\mathbf{P}_{c,:}^{\prime},\mathbf{P}_{c,:},\mathbf{P}_{c,:}).$ (5)
@@ 4 Experiment
Baselines. For long-term forecasting, we select a diverse set of state-of-the-art baselines representative of recent advancements in time series forecasting, including the Transformer-based PDF (Dai et al., 2024), the CNN-based ModernTCN (Donghao & Xue, 2024), the MLP-based TimeMixer (Wang et al., 2024), as well as other competitive methods such as iTransformer (Liu et al., 2024a), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), FEDformer (Zhou et al., 2022), MICN (Wang et al., 2022), TimesNet (Wu et al., 2023), and DLinear (Zeng et al., 2023). For short-term forecasting, we add a well-performing baseline SCINet (Liu et al., 2022a). For financial forecasting, we also incorporate the momentum strategy CSM (Jegadeesh & Titman, 1993) and the reversal strategy BLSW (Poterba & Summers, 1988), along with two classic deep learning models, LSTM (Hochreiter & Schmidhuber, 1997) and Transformer (Vaswani et al., 2017a), to provide a comprehensive evaluation.
Implementation Details. All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on two NVIDIA RTX 3090 24GB GPUs. We use the Adam optimizer (Kingma, 2014) with a learning rate selected from $\{1e\text{-}3,1e\text{-}4,5e\text{-}4\}$. The number of patches $N$ is set to 30. For additional details on hyperparameters and settings, please refer to the Appendix D.
@@ 4.1 Long-term Forecasting
Setups. We conduct long-term forecasting experiments on eight widely-used real-world datasets, including the Electricity Transformer Temperature (ETT) dataset with its four subsets (ETTh1, ETTh2, ETTm1, ETTm2), as well as Weather, Electricity, Traffic, and Solar (Liu et al., 2024a). These datasets exhibit strong non-stationary characteristics, detailed in Appendix C. Following previous works (Zhou et al., 2021; Wu et al., 2021), we use Mean Square Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. We set the input length $I$ to 720 for our method. For other baselines, we adopt two settings: one uses the original results from the baseline papers, and the other involves searching for the optimal input length $I$ and other hyperparameters. Details of the search process can be found in Sec. E.1.
Results. As shown in Tab. 1, TimeBridge consistently outperforms other methods on these non-stationary datasets, with an average improvement of over 10% compared to the baselines. Moreover, in experiments with different hyperparameter search settings in Tab. 2, TimeBridge continues to achieve the best overall results. Specifically, compared to the current state-of-the-art methods—Transformer-based PDF (Dai et al., 2024), CNN-based ModernTCN (Donghao & Xue, 2024), and MLP-based TimeMixer (Wang et al., 2024)—TimeBridge reduces MSE and MAE by 3.10%/1.64%, 3.55%/0.81%, and 6.92%/4.54%, respectively.
Models TimeBridge iTransformer PDF TimeMixer PatchTST Crossformer FEDformer ModernTCN MICN TimesNet DLinear
@@
ETTm1 0.353 0.388 0.407 0.410 0.339 0.372 0.381 0.395 0.353 0.382 0.431 0.443 0.448 0.452 0.351 0.381 0.392 0.414 0.400 0.406 0.357 0.379
ETTm2 0.246 0.310 0.288 0.332 0.252 0.313 0.275 0.323 0.256 0.317 0.621 0.510 0.305 0.349 0.253 0.314 0.290 0.343 0.291 0.333 0.267 0.332
ETTh1 0.399 0.420 0.454 0.447 0.399 0.419 0.447 0.440 0.413 0.434 0.446 0.464 0.440 0.460 0.404 0.420 0.440 0.462 0.458 0.450 0.423 0.437
ETTh2 0.346 0.394 0.383 0.407 0.327 0.376 0.364 0.395 0.324 0.381 0.835 0.675 0.434 0.447 0.322 0.379 0.411 0.440 0.414 0.427 0.431 0.447
Weather 0.218 0.259 0.258 0.279 0.225 0.261 0.240 0.271 0.226 0.264 0.343 0.386 0.309 0.360 0.224 0.264 0.243 0.299 0.259 0.287 0.246 0.300
Electricity 0.149 0.246 0.178 0.270 0.159 0.250 0.182 0.272 0.159 0.253 0.293 0.351 0.205 0.315 0.156 0.253 0.187 0.295 0.192 0.295 0.166 0.264
Traffic 0.360 0.255 0.428 0.282 0.383 0.254 0.484 0.297 0.391 0.264 0.535 0.300 0.573 0.347 0.396 0.270 0.542 0.316 0.620 0.336 0.434 0.295
Solar 0.181 0.239 0.233 0.262 0.205 0.265 0.216 0.280 0.207 0.294 0.204 0.248 0.296 0.407 0.228 0.282 0.247 0.296 0.319 0.348 0.329 0.400
  • $\dagger$
    For papers that do not originally report results on the Solar dataset, we run experiments using their official code with default settings.
Table 1: Long-term forecasting results from the original papers. All results are averaged across four different prediction lengths: $O\in\{96,192,336,720\}$. The best and second-best results are highlighted in bold and underlined, respectively. See Tab. 10 for full results.
Models TimeBridge iTransformer PDF TimeMixer PatchTST Crossformer FEDformer ModernTCN MICN TimesNet DLinear
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1 0.353 0.388 0.362 0.391 0.335 0.373 0.348 0.375 0.353 0.382 0.420 0.435 0.382 0.422 0.351 0.381 0.383 0.406 0.400 0.406 0.357 0.379
ETTm2 0.246 0.310 0.269 0.329 0.247 0.311 0.256 0.315 0.256 0.317 0.518 0.501 0.292 0.343 0.253 0.314 0.277 0.336 0.291 0.333 0.267 0.332
ETTh1 0.399 0.420 0.439 0.448 0.395 0.420 0.411 0.423 0.413 0.434 0.440 0.463 0.428 0.454 0.404 0.420 0.433 0.462 0.458 0.450 0.423 0.437
ETTh2 0.346 0.394 0.374 0.406 0.326 0.376 0.316 0.384 0.324 0.381 0.809 0.658 0.388 0.434 0.322 0.379 0.385 0.430 0.414 0.427 0.431 0.447
Weather 0.218 0.259 0.233 0.271 0.220 0.259 0.222 0.262 0.226 0.264 0.228 0.287 0.305 0.287 0.224 0.264 0.242 0.298 0.259 0.287 0.240 0.300
Electricity 0.149 0.246 0.164 0.261 0.156 0.250 0.156 0.246 0.159 0.253 0.181 0.279 0.205 0.315 0.156 0.253 0.182 0.292 0.192 0.295 0.166 0.264
Traffic 0.360 0.255 0.397 0.282 0.377 0.256 0.387 0.262 0.391 0.264 0.523 0.284 0.573 0.347 0.396 0.270 0.535 0.312 0.620 0.336 0.434 0.295
Solar 0.181 0.239 0.200 0.260 0.205 0.265 0.192 0.244 0.194 0.245 0.191 0.242 0.243 0.350 0.228 0.282 0.213 0.266 0.244 0.334 0.247 0.309
@@ 4.2 Short-term Forecasting
Setups. For short-term forecasting, we conduct experiments on the PeMS datasets (Wang et al., 2024), which capture complex spatiotemporal correlations among multiple variates across city-wide traffic networks. We use mean absolute error (MAE), mean absolute percentage error (MAPE), and root mean squared error (RMSE) as evaluation metrics. The input length $I$ is set to 96 and the output length $O$ to 12 for all baselines. Details of datasets and metrics are in Appendix C and Sec. B.2.
Results. As shown in Tab. 3, methods that perform well in long-term forecasting with channel-independent approaches, such as PatchTST (Nie et al., 2023) and DLinear (Zeng et al., 2023), suffer from significant performance degradation on the PeMS dataset due to its strong inter-variable dependencies. In contrast, TimeBridge demonstrates robust performance on this challenging task, outperforming even the recent state-of-the-art method TimeMixer (Wang et al., 2024), which highlights its effectiveness in capturing complex spatiotemporal relationships.
@@
4.3 Financial Forecasting
Setups. We conduct experiments on both the U.S. and Chinese stock markets, including the S&P 500 and CSI 500 indices. Stock price movements are influenced by various factors such as economic indicators, market sentiment, geopolitical events, and company-specific news, leading to high non-stationarity. We predict next-day returns using historical data and generate investment portfolios with a buy-hold-sell strategy (Sanderson & Lumpkin-Sowers, 2018). At day $t+1$ open, traders sell day $t$ stocks and buy top-ranked ones based on predicted returns. Following previous work (Lin et al., 2021; Li et al., 2024), we evaluate performance using Annual Return Ratio (ARR), Annual Volatility (AVol), Maximum Drawdown (MDD), Annual Sharpe Ratio (ASR), Calmar Ratio (CR), and Information Ratio (IR). Details of datasets and metrics are in Appendix C and Sec. B.3.
Results. As shown in Tab. 4, due to the non-stationary dynamics and complex market dependencies, other baseline methods struggle to consistently and accurately identify broadly optimal portfolios across different markets. In contrast, TimeBridge consistently performs best in both financial markets, demonstrating its ability to capture short-term fluctuations within financial time series and long-term cointegration between sectors.

### v3

@@ 1 Introduction
Figure 1: Visualization of the impact of non-stationarity on short-term and long-term modeling. The goal is to forecast two cointegrated sequences, $X_{t}$ and $Y_{t}$, where $X_{t}$ exhibits two random fluctuations. (a) Retaining non-stationarity preserves long-term cointegration between variates but leads to spurious regressions in short-term modeling (orange line). (b) Removing non-stationarity avoids short-term spurious regressions but disrupts long-term similar trends (green line).
Recently, many methods have emerged to utilize a normalization-and-denormalization paradigm to address non-stationarity in time series (Kim et al., 2022; Fan et al., 2023; Liu et al., 2023, 2024b). For instance, RevIN (Kim et al., 2022) normalizes the input data and subsequently applies its distributional characteristics to denormalize the output predictions. Building on this approach, other methods have designed distributional prediction networks (Fan et al., 2023) and more refined normalization techniques (Liu et al., 2023) to further mitigate non-stationarity. On the other hand, some studies (Liu et al., 2022b; Ma et al., 2024; Fan et al., 2024) argue that over-stabilizing time series may actually reduce the richness of embedded features, leading to a decline in model performance. Existing methods for addressing non-stationarity in time series face a dilemma: some prioritize eliminating non-stationary factors to reduce overfitting, while others attempt to incorporate these factors but lack comprehensive theoretical frameworks. Furthermore, they do not adequately explain the trade-offs between removing non-stationarity and leveraging it for modeling.
It is notable that non-stationary characteristics have distinct impacts on modeling short-term and long-term dependencies. Non-stationarity can lead to spurious regressions in short-term modeling due to the high randomness and unpredictability of short-term fluctuations (Noriega & Ventosa-Santaulària, 2007) (see Figure 1a). For example, a sudden drop in temperature could be caused by a typhoon or cold front, both of which have no intrinsic connection. Retaining non-stationarity can result in false correlations between such unrelated events. However, non-stationarity is crucial for capturing long-term cointegration relationships among variates, reflecting their co-movement or synchronized changes over time (Fanchon & Wendel, 1992). Removing non-stationarity may also eliminate these essential long-term dependencies (see Figure 1b). Therefore, while non-stationarity can cause spurious regressions in short-term modeling, it is essential for modeling long-term dependencies between variates. Conversely, eliminating non-stationarity benefits short-term modeling but erases long-term cointegration.
In this paper, to address the dual challenges posed by non-stationarity in short-term and long-term modeling, we propose distinct strategies tailored to each scenario. For short-term modeling, we eliminate non-stationarity to capture the strong temporal dependencies within each variate, as short-term causal relationships mainly exist between consecutive time points within a single variate rather than across variates. This strategy reduces the risk of spurious regressions from non-stationary fluctuations, enabling the model to better capture the local causal dynamics. For long-term modeling, we utilize the preserved non-stationarity to uncover long-term cointegration relationships between different variates, thereby enabling more accurate and reliable long-term forecasting.
@@ 2 Related Works
As shown in Figure 2, recent advancements in multivariate time series forecasting have predominantly focused on two core directions: Normalization and Dependecy Modeling.
Normalization can be divided into stationary and non-stationary methods. Stationary methods (Kim et al., 2022; Fan et al., 2023; Liu et al., 2024b, 2023) aim to eliminate non-stationarity through model-agnostic normalization techniques, thereby preventing spurious regressions and enhancing model performance. For example, RevIN (Kim et al., 2022) applies Z-normalization to the input sequence and then reverses the normalization on the output using the distributional characteristics of the input, assuming that both share similar distributional properties. Dish-TS (Fan et al., 2023) takes this further by predicting the statistical characteristics of the output with a distribution prediction model. Additionally, SAN (Liu et al., 2023) offers a more granular patch-level prediction method. Conversely, some approaches (Liu et al., 2022b; Ma et al., 2024; Fan et al., 2024) advocate preserving non-stationarity, as excessive normalization can eliminate inherent diverse sequence characteristics and limit predictive accuracy.
@@ 3.1 Structure Overview
As illustrated in Figure 3, our proposed TimeBridge consists of four key components: (a) Patch Embedding segments the input sequence into non-overlapping patches and transforms each patch into a patch token; (b) Integrated Attention models the dependencies among all patch tokens of the same variates. By eliminating non-stationarity within each patch token, it mitigates the risk of spurious regressions that could arise from abrupt short-term changes; (c) Patch Downsampling aggregates global information and reduces the number of patches to encapsulate richer long-term features within each patch, while simultaneously lowering computational complexity; (d) Cointegrated Attention preserves the non-stationary characteristics of the sequence and models the long-term cointegration relationships across different variates within the same temporal window.
@@ 3.2 Patch Embedding
$\{\mathbf{p}_{1},\cdots,\mathbf{p}_{N}\}=\text{Patching}(\mathbf{X}),$ (1)
$\mathbf{P}=\text{Embedding}(\mathbf{p}_{1},\cdots,\mathbf{p}_{N})$ (2)
@@ 3.3 Integrated Attention
$\mathbf{p}_{i}^{\text{Trend}}=\text{AvgPool}$ $(\text{Padding}(\mathbf{p}_{i})),\mathbf{p}_{i}^{\prime}=\mathbf{p}_{i}-\mathbf{p}_{i}^{\text{Trend}}$ (3)
$\mathbf{P}_{c,:}^{\prime}$ =$=\{\mathbf{p}_{1}^{\prime},\cdots,\mathbf{p}_{N}^{\prime}\},$ (4)
@@ 3.4 Patch Downsampling
$\mathbf{P}_{c,:}^{\prime}$ $=\text{Downsample}(\mathbf{P}_{c,:})$ (7)
$\mathbf{P}_{c,:}$ $=\text{Attention}(\mathbf{P}_{c,:}^{\prime},\mathbf{P}_{c,:},\mathbf{P}_{c,:}).$ (8)
@@ 4 Experiment
Baselines. For long-term forecasting, we select a diverse set of state-of-the-art baselines representative of recent advancements in time series forecasting, including the Transformer-based DeformableTST (Luo & Wang, 2024), the CNN-based ModernTCN (Donghao & Xue, 2024), the MLP-based TimeMixer (Wang et al., 2024b), as well as other competitive methods such as iTransformer (Liu et al., 2024a), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), FEDformer (Zhou et al., 2022), MICN (Wang et al., 2022), TimesNet (Wu et al., 2023), and DLinear (Zeng et al., 2023). For short-term forecasting, we add a well-performing baseline SCINet (Liu et al., 2022a). For financial forecasting, we also incorporate the momentum strategy CSM (Jegadeesh & Titman, 1993) and the reversal strategy BLSW (Poterba & Summers, 1988), along with two classic deep learning models, LSTM (Hochreiter & Schmidhuber, 1997) and Transformer (Vaswani et al., 2017a), to provide a comprehensive evaluation.
Implementation Details. All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on two NVIDIA RTX 3090 24GB GPUs. We use the Adam optimizer (Kingma, 2014) with a learning rate selected from $\{1e\text{-}3,1e\text{-}4,5e\text{-}4\}$. The number of patches $N$ is set to 30. We adopt a hybrid MAE loss that operates in both the time and frequency domains for stable training (Wang et al., 2024a). For additional details on hyperparameter settings and loss function, please refer to the Appendix D.
@@ 4.1 Long-term Forecasting
Setups. We conduct long-term forecasting experiments on several widely-used real-world datasets, including the Electricity Transformer Temperature (ETT) dataset with its four subsets (ETTh1, ETTh2, ETTm1, ETTm2) (Wu et al., 2021), as well as Weather, Electricity, Traffic, Solar (Liu et al., 2024a) and Climate (Godahewa et al., 2021). These datasets exhibit strong non-stationary characteristics, detailed in Appendix C. Following previous works (Zhou et al., 2021; Wu et al., 2021), we use Mean Square Error (MSE) and Mean Absolute Error (MAE) as evaluation metrics. We set the input length $I$ to 720 for our method. For other baselines, we adopt the setting that searches for the optimal input length $I$ and other hyperparameters. Details of the metric and the searching process can be found in Sec. B.1 and Sec. E.1.
Results. As shown in Tab. 1, TimeBridge consistently achieves the best overall performance. Notably, the large-scale Climate dataset, with its 1763 channels, presents substantial challenges due to its high dimensionality and intricate temporal dependencies. TimeBridge adeptly balances non-stationarity and dependency modeling, achieving consistently strong predictive performance. Quantitatively, compared to state-of-the-art methods—Transformer-based DeformableTST (Luo & Wang, 2024), CNN-based ModernTCN (Donghao & Xue, 2024), and MLP-based TimeMixer (Wang et al., 2024b)—TimeBridge reduces MSE and MAE by $1.85\%/2.49\%$, $5.56\%/4.12\%$, and $13.66\%/7.58\%$, respectively.
Models TimeBridge iTransformer
DeformableTST
TimeMixer PatchTST Crossformer FEDformer ModernTCN MICN TimesNet DLinear
@@
ETTm1 0.344 0.379 0.362 0.391 0.348 0.383 0.355 0.380 0.353 0.382 0.420 0.435 0.382 0.422 0.351 0.381 0.383 0.406 0.400 0.406 0.357 0.379
ETTm2 0.246 0.310 0.269 0.329 0.257 0.319 0.257 0.318 0.256 0.317 0.518 0.501 0.292 0.343 0.253 0.314 0.277 0.336 0.291 0.333 0.267 0.332
ETTh1 0.397 0.424 0.439 0.448 0.404 0.423 0.427 0.441 0.413 0.434 0.440 0.463 0.428 0.454 0.404 0.420 0.433 0.462 0.458 0.450 0.423 0.437
ETTh2 0.341 0.382 0.374 0.406 0.328 0.377 0.349 0.397 0.324 0.381 0.809 0.658 0.388 0.434 0.322 0.379 0.385 0.430 0.414 0.427 0.431 0.447
Weather 0.219 0.249 0.233 0.271 0.222 0.262 0.226 0.264 0.226 0.264 0.228 0.287 0.305 0.287 0.224 0.264 0.242 0.298 0.259 0.287 0.240 0.300
Electricity 0.149 0.245 0.164 0.261 0.161 0.261 0.185 0.284 0.159 0.253 0.181 0.279 0.205 0.315 0.156 0.253 0.182 0.292 0.192 0.295 0.166 0.264
Traffic 0.360 0.255 0.397 0.282 0.391 0.278 0.409 0.279 0.391 0.264 0.523 0.284 0.573 0.347 0.396 0.270 0.535 0.312 0.620 0.336 0.434 0.295
Solar 0.181 0.239 0.200 0.260 0.185 0.254 0.193 0.252 0.194 0.245 0.191 0.242 0.243 0.350 0.228 0.282 0.213 0.266 0.244 0.334 0.247 0.309
Climate 1.057 0.494 1.060 0.496 1.060 0.496 1.414 0.606 1.387 0.562 1.170 0.550 1.378 0.662 1.155 0.542 1.166 0.570 1.295 0.569 1.134 0.541
@@ 4.2 Short-term Forecasting
Setups. For short-term forecasting, we conduct experiments on the PeMS datasets (Wang et al., 2024b), which capture complex spatiotemporal correlations among multiple variates across city-wide traffic networks. We use mean absolute error (MAE), mean absolute percentage error (MAPE), and root mean squared error (RMSE) as evaluation metrics. The input length $I$ is set to 96 and the output length $O$ to 12 for all baselines. Details of datasets and metrics are in Appendix C and Sec. B.2.
Results. As shown in Tab. 2, methods that perform well in long-term forecasting with channel-independent approaches, such as PatchTST (Nie et al., 2023) and DLinear (Zeng et al., 2023), suffer from significant performance degradation on the PeMS dataset due to its strong inter-variable dependencies. In contrast, TimeBridge demonstrates robust performance on this challenging task, outperforming even the recent state-of-the-art method TimeMixer (Wang et al., 2024b), which highlights its effectiveness in capturing complex spatiotemporal relationships.
@@
4.3 Financial Forecasting
Setups. We conduct experiments on both the U.S. and Chinese stock markets, including the S&P 500 and CSI 500 indices. Stock price movements are influenced by various factors such as economic indicators, market sentiment, geopolitical events, and company-specific news, leading to high non-stationarity. We predict next-day returns using historical data and generate investment portfolios with a buy-hold-sell strategy (Sanderson & Lumpkin-Sowers, 2018). At day $t+1$ open, traders sell day $t$ stocks and buy top-ranked ones based on predicted returns. Following previous work (Lin et al., 2021), we evaluate performance using Annual Return Ratio (ARR), Annual Volatility (AVol), Maximum Drawdown (MDD), Annual Sharpe Ratio (ASR), Calmar Ratio (CR), and Information Ratio (IR). Details of datasets and metrics are in Appendix C and Sec. B.3.
Results. As shown in Tab. 3, the inherent non-stationarity and intricate dependencies within financial markets make it difficult for baseline methods to consistently identify optimal portfolios across different markets. While financial market fluctuations are notoriously unpredictable, TimeBridge adapts well to these challenges. By capturing short-term volatility within financial time series and preserving long-term cointegration across sectors, it achieves consistently strong performance, outperforming existing methods in overall market efficiency.